A landscape architect designs a garden with a perimeter of 44 feet. The width is 2 feet less than the length. Find the length and the width.

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)


If it's a rectangle:
L + W = 22
L - W = 2
-------------- Add

Answer by greenestamps(13198)



A simple solution using logical reasoning instead of formal algebra....

Increase the width by 2 feet.
Now the garden is a square with a perimeter of 44+4 = 48 feet.
That means the new garden is a square with side length 48/4 = 12.
Now take away those 2 extra feet you added to the width.
The garden the architect designed has length 12 and width 10.